In the following one solution of a second y1 order linear homogene DE is given. Find the second linearly independent solution y2 using the method of reduction of order.




2x²y" + 3xy'-y=0 ,. y1=(1/x)

(1-x2)y"-2xy'+2y=0 , y1=x

x2y"+2xy'-2y=0,. y1=x

x2y"+3xy'+y=0, y1=1/x

x2y"-x(x+2)y'=0,. y1=x


1
Expert's answer
2021-06-28T04:00:37-0400

1.


"2x^2y'' + 3xy'-y=0 ,y_1=(1\/x)"

"y(x)=u(x)y_1(x)"

"y(x)=u(x)x^{-1}"

"y'=u'x^{-1}-ux^{-2}"

"y''=u''x^{-1}-2u'x^{-2}+2ux^{-3}"


"2x^2(u''x^{-1}-2u'x^{-2}+2ux^{-3})+3x(u'x^{-1}-ux^{-2})"

"-ux^{-1}=0"

"2xu''-4u'+4ux^{-1}+3u'-3ux^{-1}-ux^{-1}=0"

"2xu''-u'=0"

"u'=w, u''=w'"

"2xw'-w=0"

"\dfrac{dw}{w}=\dfrac{dx}{2x}"

"\int\dfrac{dw}{w}=\int\dfrac{dx}{2x}"

"\ln|w|=\dfrac{1}{2}\ln|x|+\ln C_1"

"w=C_1\sqrt{x}"

"u'=C_1\sqrt{x}"

"\int du=\int C_1\sqrt{x}dx"

"u=C_2x^{3\/2}+C_3"

"y(x)=u(x)x^{-1}"

"y(x)=C_2x^{1\/2}+C_3x^{-1}"

2.


"(1-x^2)y''-2xy'+2y=0 , y_1=x"

"y(x)=u(x)y_1(x)"

"y(x)=u(x)x"

"y'=u'x+u"

"y''=u''x+2u'"


"(1-x^2)(u''x+2u')-2x(u'x+u)+2ux=0"

"x(1-x^2)u''+2u'-2x^2u'-2x^2u'-2ux+2ux=0"

"x(1-x^2)u''+(2-4x^2)u'=0"

"u'=w, u''=w'"

"x(1-x^2)w'+(2-4x^2)w=0"

"\dfrac{dw}{w}=\dfrac{2-4x^2}{x(1-x^2)}dx"

"\int\dfrac{dw}{w}=\int\dfrac{2-4x^2}{x(1-x^2)}dx"

"\dfrac{2-4x^2}{x(1-x^2)}dx=\dfrac{A}{x}dx+\dfrac{B}{1-x}dx+\dfrac{C}{1+x}dx"

"A(1-x^2)+Bx(1+x)+Cx(1-x)=2-4x^2"

"A-Ax^2+Bx+Bx^2+Cx-Cx^2=2-4x^2"

"A=2"

"B+C=0"

"-A+B-C=-4"

"A=2, B=-1, C=1"

"\int\dfrac{dw}{w}=\int\dfrac{2}{x}dx-\int\dfrac{1}{1-x}dx+\int\dfrac{1}{1+x}dx"

"\ln|w|=2\ln|x|+\ln|1-x|+\ln|1+x|+\ln C_1"

"w=C_1x^2(1-x^2)"

"u'=C_1x^2(1-x^2)"

"u=\dfrac{1}{3}C_1x^3-\dfrac{1}{5}C_1x^5+C_2"

"y=\dfrac{1}{3}C_1x^4-\dfrac{1}{5}C_1x^6+C_2x"



3.


"x^2y''+2xy-2y=0 , y_1=x"

"y(x)=u(x)y_1(x)"

"y(x)=u(x)x"

"y'=u'x+u"

"y''=u''x+2u'"




"x^2(u''x+2u')+2x(u'x+u)-2ux=0"

"x^3u''+2x^2u'+2ux-2ux=0"

"x^3u''+2x^2u'=0"

"u'=w, u''=w'"

"x^3w'+2x^2w=0"

If "x\not=0"

"\dfrac{dw}{w}=-\dfrac{2}{x}dx"

"\int\dfrac{dw}{w}=-\int\dfrac{2}{x}dx"

"\ln|w|=-2\ln|x|+\ln C_1"

"w=\dfrac{C_1}{x^2}"

"u'=\dfrac{C_1}{x^2}"

"u=-\dfrac{C_1}{x}+C_2"

"y=C_3+C_2x"



4.


"x^2y'' + 3xy'+y=0 ,y_1=(1\/x)"

"y(x)=u(x)y_1(x)"

"y(x)=u(x)x^{-1}"

"y'=u'x^{-1}-ux^{-2}"

"y''=u''x^{-1}-2u'x^{-2}+2ux^{-3}"


"x^2(u''x^{-1}-2u'x^{-2}+2ux^{-3})+3x(u'x^{-1}-ux^{-2})"

"+ux^{-1}=0"

"xu''-2u'+2ux^{-1}+3u'-3ux^{-1}+ux^{-1}=0"

"xu''+u'=0"

"u'=w, u''=w'"

"xw'+w=0"

"\dfrac{dw}{w}=-\dfrac{dx}{x}"

"\int\dfrac{dw}{w}=-\int\dfrac{dx}{x}"

"\ln|w|=-\ln|x|+\ln C_1"

"w=\dfrac{C_1}{x}"

"u'=\dfrac{C_1}{x}"

"\int du=\int \dfrac{C_1}{x}dx"

"u=C_1\ln|x|+C_2"

"y(x)=u(x)x^{-1}"

"y(x)=\dfrac{C_1\ln x}{x}+\dfrac{C_2}{x}"

5.


"x^2y''-x(x+2)y'=0 , y_1=x"

If "x\not=0"


"xy''-(x+2)y'=0"

"y'=w, y''=w'"

"xw'-(x+2)w=0"

If "x\not=0"

"\dfrac{dw}{w}=\dfrac{x+2}{x}dx"

"\int\dfrac{dw}{w}=\int\dfrac{x+2}{x}dx"

"\ln |w|=x+\ln x^2+\ln C_1"

"y'=C_1x^2e^x"

"y=\int C_1x^2e^xdx""\int C_1x^2e^xdx=C_1x^2e^x-2C_1\int xe^xdx"

"=C_1x^2e^x-2C_1xe^x+2C_1\int e^xdx"

"=C_1x^2e^x-2C_1xe^x+2C_1e^x+C_2"

"y=C_1x^2e^x-2C_1xe^x+2C_1e^x+C_2"